Exercise 11-1 Compute the Return on Investment (ROI) [LO11-1] Alyeska Services Company, a division of a major oil company, provides various services to the operators of the North Slope oil field in Alaska. Data concerning the most recent year appear below: Sales Net operating income Average operating assets $18,900,000 $ 4,700,000 $35,900,000 Required 1. Compute the margin for Alyeska Services Company. (Round your answer to 2 decimal places.) Margin 2. Compute the turnover for Alyeska Services Company. (Round your answer to 2 decimal places.) urnover 3. Compute the return on investment (ROI) for Alyeska Services Company. (Round your intermediate calculations and final answer to 2 decimal places.) ROl
Answer #1

1.Margin=Net operating income/Sales

=(4,700,000/18,900,000)=24.87%(Approx).

2.Turnover=Sales/Average operating assets

=(18,900,000/35,900,000)=0.53(Approx)

3.ROI=Net operating income/Average operating assets

=(4,700,000/35,900,000)=13.09%(Approx).